OpenStudy (anonymous):

A new car worth 21,000 is depreciating in value by $3,000 per year. After how many years will the car value be $ 3,000?

OpenStudy (anonymous):

21,000 - 3,000y = 3,000 -> y = 18/3 = 6 yrs